Logiweb bibliography of sup

Up Help

[0] sup

[1] base

[2] check

[3] kvanti

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-12-29.UTC:09:42:35.018035 = MJD-54098.TAI:09:43:08.018035 = LGT-4674102188018035e-6